# “Your Doctor is Spying on You”: An Analysis of Data Practices in Mobile Healthcare Applications

### v1

@@ I Introduction
Mobile healthcare (mHealth) applications are reshaping digital health by enabling continuous, remote patient–provider interaction across diagnostics, treatment, and monitoring workflows. They deliver capabilities ranging from real-time vital sign monitoring and medication adherence tracking to teleconsultations and insurance claims management [1, 2]. However, their deep integration with mobile devices and direct handling of Protected Health Information (PHI) greatly expands the attack surface for privacy breaches and malicious exploitation [3, 4, 5, 6, 7, 8]. Breaches of PHI can cause severe and long-lasting harm, including large-scale identity theft, fraudulent insurance claims, targeted discrimination, and psychological distress [9, 10, 11, 12]. Regulatory frameworks such as HIPAA [13] and GDPR [14] mandate strict controls over data collection, storage, transmission, and sharing. However, enforcement remains inconsistent, particularly for consumer-oriented mHealth applications that fall outside medical device classifications [15, 16, 17, 18, 19, 20]. This regulatory blind spot enables insecure default configurations, deprecated cryptographic protocols, and opaque data-sharing practices to proliferate in app marketplaces with minimal oversight.
@@ II Background
Security and privacy in mobile applications are critical, particularly in open ecosystems like Android, where platforms such as Google Play allow unrestricted app publication [21, 5, 22, 23, 24, 25, 26]. While this openness fosters innovation, it also introduces risks of security flaws, privacy breaches, and regulatory noncompliance [27]. These risks are amplified in healthcare apps that handle PHI and are subject to strict legal and ethical standards. Pham et al. [28] investigated privacy issues in mHealth apps but analyzed a limited sample and overlooked user experience aspects. Other studies have examined runtime behaviors, usability, and permission misuse [29, 27], yet none have simultaneously considered undeclared permissions, cryptographic weaknesses, and user concerns in healthcare. This study addresses that gap.
@@ II-A Tools
Over the past decade, open-source and academic communities have developed mature frameworks for static and dynamic mobile app analysis, enabling the detection of insecure permissions, weak cryptography, and privacy-invasive behaviors beyond superficial inspection. We employ three complementary tools in this study. The Mobile Security Framework (MobSF)[30] performs static APK analysis to identify vulnerabilities such as weak hash functions, improper intent handling, cleartext traffic, and excessive permission requests, assigning severity scores for comparative profiling[31, 32, 33, 34, 35, 36]. RiskInDroid[37] extends permission analysis by categorizing permissions into declared, used, undeclared but used, and required but unused [38]. This approach is effective for detecting stealthy access to sensitive APIs in healthcare contexts. The OWASP Mobile Audit[39], aligned with the OWASP Mobile Application Security Testing Guide (MASTG)[40], applies Static Application Security Testing (SAST) to uncover vulnerabilities ranked from critical to low severity [41]. We also reference auxiliary tools such as APKLeaks[42] and AndroBugs [43] for lightweight scanning, and note that hybrid fuzzing and penetration testing [44] expand methodological coverage. These capabilities are integrated into a reproducible pipeline tailored for mHealth, targeting code-level threats and privacy practices that circumvent platform enforcement.
@@ II-B Mobile Application Privacy and Security
Mobile applications frequently request access to sensitive data and services beyond what is functionally necessary [45, 46, 47, 48, 49, 50, 51]. While Android’s permission model is designed to promote transparency, users often approve requests without understanding their implications [52, 53, 54, 55]. As a result, permissions may be legally obtained but ethically questionable, particularly when purposes are obscured or undeclared. Prior studies [56, 57, 58] report persistent requests for access to location, camera, SMS, and contact data without clear justification. Liu et al.[59] found that developers frequently violate the principle of least privilege, requesting access to sensitive resources without functional necessity. Our findings confirm and extend these trends in the mHealth domain. Zhang et al.[60] identified dark patterns, which are manipulative UI designs that nudge users into granting data access. Such practices reduce transparency and exploit cognitive biases [61]. Tene et al.[62] advocate enforceable, user-centric privacy frameworks, yet enforcement under GDPR and HIPAA remains inconsistent, particularly for consumer-facing health apps that are not subject to medical device regulation[63]. Our findings highlight this gap between regulatory intent and practical enforcement, demonstrating that even highly popular healthcare apps bypass platform protections and fail to meet privacy-by-design standards.
@@ II-C Healthcare Applications
mHealth applications manage PHI such as clinical histories, biometric readings, medication records, and behavioral data, making data confidentiality and integrity both technical and ethical imperatives. Mosa et al.[64] classified mHealth into domains ranging from disease management to patient education, revealing highly variable risk profiles. This diversity complicates the uniform application of security standards. Regulatory fragmentation exacerbates the issue. In the United States, only apps that replace or augment clinical devices fall under HIPAA and FDA oversight[59], leaving wellness and lifestyle apps largely unregulated [65]. Reddy et al.[66] documented major breaches linked to insecure healthcare apps, often caused by outdated encryption, absent access controls, or weak transport-layer security. Park et al.[67] and Rafique et al. [68] similarly report persistent use of deprecated algorithms, insecure data storage, and excessive permission requests. An additional risk arises from data brokerage. Shah et al.[69] found that many mHealth apps embed advertising SDKs and analytics services that exfiltrate user data without meaningful consent, effectively monetizing sensitive health information[70]. While GDPR and HIPAA establish strong principles, enforcement remains inconsistent. Liu et al.[59] call for extending regulation to all apps handling PHI, regardless of classification. Amin et al.[71] explore how to balance privacy preservation with secure data handling in compliance contexts. Our study addresses these blind spots through a fine-grained, tool-assisted analysis of $272$ Android healthcare apps, identifying permission abuses, reliance on legacy cryptography, and insecure certificate handling. By triangulating technical findings with over 2.52.5M user reviews, we reveal a consistent pattern in which usability failures and opaque data practices undermine trust in digital healthcare.
@@ IV-D Review Analysis
We conducted a large-scale sentiment and term-frequency analysis on user feedback for all $272$ applications. Using a Google Play Store scraper, we collected more than 2.52.5M reviews and processed them with the TextBlob NLP library for sentiment classification. To reduce noise from bots or low-information feedback, we excluded all reviews with fewer than $100$ characters. After filtering, each application averaged $8,781$ reviews: $6,309$ positive, $1,451$ negative, and $1,020$ neutral. To map user-reported concerns to technical domains, we developed five semantic categories: confidentiality, user information, permissions, account access, and usable security. Term extraction and classification revealed that usable security issues dominated, appearing in $553,495$ reviews. These complaints, consistent with prior observations by Diao et al. [72], referenced interface instability, slow performance, non-responsive features, and repeated authentication failures. Account access issues appeared in $154,032$ reviews, while permissions and user information were mentioned in $38,126$ and $38,883$ reviews respectively. Confidentiality concerns, including explicit mentions of privacy violations and PHI exposure, occurred in $15,176$ reviews. The category distributions are summarised in Table III. The results show that usability failures in authentication and data access closely align with security weaknesses, indicating shared engineering flaws that undermine both stability and user trust.
@@ VI-A Platform Policy Enforcement
In this study, we examined Android applications, whose permission framework is designed for least-privilege use, yet enforcement is uneven. Many apps in our dataset bypass consent by exploiting undeclared or proprietary permissions, often accessing sensitive data such as location, SMS, and device identifiers. Google Play’s review process prioritizes malware detection but rarely flags static privacy violations, including insecure certificate configurations and privileged API use [73, 74]. To close this gap, marketplaces should embed automated tools such as RiskInDroid [37], MobSF [30], and OWASP Mobile Audit [39] directly into submission workflows. Extending these checks to silent permission use, legacy cryptography, and telephony misuse, combined with manifest-to-behavior audits and update-time differential analysis, would help detect and prevent privilege escalation.
@@ VI-B By-Design Engineering
Weak enforcement is compounded by insecure development practices. Nearly half of evaluated apps still rely on deprecated cryptography (SHA-1, MD5), and known exploits such as Janus and StrandHogg 2.0 remain exploitable. This reflects inconsistent integration of SAST/DAST tools into CI/CD pipelines. Developers should adopt the OWASP MASVS [40], embed security linters into IDEs, and align authentication with NIST SP 800-63B [75]. To address risks from third-party SDKs, SBOM generation and automated dependency scanning should become prerequisites for app store submission [76].
Users notice when app behavior conflicts with stated purpose. Permission–function mismatches, especially for location, SMS, and contacts, frequently erode trust [77]. Progressive, context-aware permission flows [78] can reduce suspicion, while platform rules should prohibit dark patterns that coerce consent, in line with EU Digital Services Act requirements [79]. At scale, NLP-based monitoring of app reviews could surface privacy concerns early, allowing developers to intervene before reputational damage occurs.
@@ VI-C Regulation of Gray-Zone Health Apps
Trust also depends on regulatory coverage. Frameworks such as HIPAA and GDPR focus on clinical providers, leaving many gray-zone apps (e.g., fitness trackers, symptom checkers) outside formal oversight despite their collection of PHI-equivalent data [80]. Regulators should broaden the definition of health data handlers to cover any entity processing biometric or medical information [63]. Privacy nutrition labels in app stores and independent certification schemes (e.g., SOC 2–style for mHealth) would provide clearer accountability and verifiable standards [81].
@@ VI-D Automated Vetting and Socio-Technical Oversight
Finally, sustaining improvements requires moving beyond manual audits. Continuous vetting pipelines that combine static and dynamic analysis, taint tracking, permission mapping, and traffic anomaly detection can expose covert channels and privilege escalation paths [82]. Sandboxed monitoring of CPU, sensor, and network activity adds runtime assurance. For decision-makers, explainable dashboards built on tools like MobSF and RiskInDroid can translate technical findings into actionable insights [76], supporting coordinated socio-technical oversight.

### v2

@@ I Introduction
Mobile healthcare (mHealth) applications are reshaping digital health by enabling continuous, remote patient–provider interaction across diagnostics, treatment, and monitoring workflows. They deliver capabilities ranging from real-time vital sign monitoring and medication adherence tracking to teleconsultations and insurance claims management [mosa2012systematic, 4278593]. However, their deep integration with mobile devices and direct handling of Protected Health Information (PHI) greatly expands the attack surface for privacy breaches and malicious exploitation [duckert2022protecting, tazi2024evaluating, saka2024evaluating, das2022privacy, tazi2024sok, tazi2022sok]. Breaches of PHI can cause severe and long-lasting harm, including large-scale identity theft, fraudulent insurance claims, targeted discrimination, and psychological distress [6906856, osawaru2024electronic, surani2023security, surani2022understanding]. Regulatory frameworks such as HIPAA [6845198] and GDPR [tzanou2023health] mandate strict controls over data collection, storage, transmission, and sharing. However, enforcement remains inconsistent, particularly for consumer-oriented mHealth applications that fall outside medical device classifications [grover2025sok, adhikari2025natural, tazi2024we, adhikari2023evolution, tazi2024improving, monroe2021location]. This regulatory blind spot enables insecure default configurations, deprecated cryptographic protocols, and opaque data-sharing practices to proliferate in app marketplaces with minimal oversight.
@@ II Background
Security and privacy in mobile applications are critical, particularly in open ecosystems like Android, where platforms such as Google Play allow unrestricted app publication [das2024design, saka2024evaluating, kishnani2023assessing, kishnani2024towards, kishnani2024dual, kishnani2022privacy, neupane2022data]. While this openness fosters innovation, it also introduces risks of security flaws, privacy breaches, and regulatory noncompliance [taylor2016quantifying]. These risks are amplified in healthcare apps that handle PHI and are subject to strict legal and ethical standards. Sun et al. [sun2020oat] studied attack detection and integrity in healthcare embedded systems. Other studies have examined runtime behaviors, usability, and permission misuse [marky2020don, taylor2016quantifying], yet none have simultaneously considered undeclared permissions, cryptographic weaknesses, and user concerns in healthcare. This study addresses that gap.
@@ II-A Tools
Over the past decade, open-source and academic communities have developed mature frameworks for static and dynamic mobile app analysis, enabling the detection of insecure permissions, weak cryptography, and privacy-invasive behaviors beyond superficial inspection. We employ three complementary tools in this study. The Mobile Security Framework (MobSF) [bergadano2020modular] performs static APK analysis to identify vulnerabilities such as weak hash functions, improper intent handling, cleartext traffic, and excessive permission requests, assigning severity scores for comparative profiling [tileria2020wearflow, kalhor2023evaluating, jensen2021multi, noah2024aging, noah2024evaluating, noah2022privacy]. RiskInDroid [merlo2017riskindroid] extends permission analysis by categorizing permissions into declared, used, undeclared but used, and required but unused [khorkheli2024improving]. This approach is effective for detecting stealthy access to sensitive APIs in healthcare contexts. The OWASP Mobile Audit[acharya2015owasp], aligned with the OWASP Mobile Application Security Testing Guide (MASTG)[holguera2022owasp], applies Static Application Security Testing (SAST) to uncover vulnerabilities ranked from critical to low severity [kishnani2024securing]. We also reference auxiliary tools such as APKLeaks [apkleaks2021] and AndroBugs [AndroBugs2023framework] for lightweight scanning, and note that hybrid fuzzing and penetration testing [pandey2021web] expand methodological coverage. These capabilities are integrated into a reproducible pipeline tailored for mHealth, targeting code-level threats and privacy practices that circumvent platform enforcement.
@@ II-B Mobile Application Privacy and Security
Mobile applications frequently request access to sensitive data and services beyond what is functionally necessary [leontiadis2012don, momenzadeh2020bayesian, das2020user, tazi2023accessibility, sun2016blender, momenzadeh2021bayesian, podapati2025sok]. While Android’s permission model is designed to promote transparency, users often approve requests without understanding their implications [fife2012privacy, streiff2019overpowered, debnath2020studies, dev2019personalized]. As a result, permissions may be legally obtained but ethically questionable, particularly when purposes are obscured or undeclared. Prior studies [10.1145/2420950.2420956, tvidas2011all, sivan2019analysis] report persistent requests for access to location, camera, SMS, and contact data without clear justification. Liu et al. [LIU20112022] found that developers frequently violate the principle of least privilege, requesting access to sensitive resources without functional necessity. Our findings confirm and extend these trends in the mHealth domain. Zhu et al. [10.1145/2623330.2623705] advocate enforceable, user-centric privacy frameworks, yet enforcement under GDPR and HIPAA remains inconsistent, particularly for consumer-facing health apps that are not subject to medical device regulation [carayon2018challenges]. Our findings highlight this gap between regulatory intent and practical enforcement, demonstrating that even highly popular healthcare apps bypass platform protections and fail to meet privacy-by-design standards.
@@ II-C Healthcare Applications
mHealth applications manage PHI such as clinical histories, biometric readings, medication records, and behavioral data, making data confidentiality and integrity both technical and ethical imperatives. Prasanna et al.[prasanna2013decision] classified mHealth into domains ranging from disease management to patient education, revealing highly variable risk profiles. This diversity complicates the uniform application of security standards. Regulatory fragmentation exacerbates the issue. In the United States, only apps that replace or augment clinical devices fall under HIPAA and FDA oversight[LIU20112022], leaving wellness and lifestyle apps largely unregulated [moen2005health]. Reddy et al. [reddy2023review] documented major breaches linked to insecure healthcare apps, often caused by outdated encryption, absent access controls, or weak transport-layer security. Park and Youm [park2022appsecurity] and Ahmad et al. [ahmad2023leveraging] similarly report persistent use of deprecated algorithms, insecure data storage, and excessive permission requests. An additional risk arises from data brokerage. Shah and Lim [shah2024analysis] found that many mHealth apps embed advertising SDKs and analytics services that exfiltrate user data without meaningful consent, effectively monetizing sensitive health information[fernandez2018analytic]. While GDPR and HIPAA establish strong principles, enforcement remains inconsistent. Liu et al.[LIU20112022] call for extending regulation to all apps handling PHI, regardless of classification. Amin et al.[amin2024balancing] explore how to balance privacy preservation with secure data handling in compliance contexts. Our study addresses these blind spots through a fine-grained, tool-assisted analysis of $272$ Android healthcare apps, identifying permission abuses, reliance on legacy cryptography, and insecure certificate handling. By triangulating technical findings with over 2.52.5M user reviews, we reveal a consistent pattern in which usability failures and opaque data practices undermine trust in digital healthcare.
@@ IV-D Review Analysis
We conducted a large-scale sentiment and term-frequency analysis on user feedback for all $272$ applications. Using a Google Play Store scraper, we collected more than 2.52.5M reviews and processed them with the TextBlob NLP library for sentiment classification. To reduce noise from bots or low-information feedback, we excluded all reviews with fewer than $100$ characters. After filtering, each application averaged $8,781$ reviews: $6,309$ positive, $1,451$ negative, and $1,020$ neutral. To map user-reported concerns to technical domains, we developed five semantic categories: confidentiality, user information, permissions, account access, and usable security. Term extraction and classification revealed that usable security issues dominated, appearing in $553,495$ reviews. These complaints, consistent with prior observations by Diao et al. [diao2019kindness], referenced interface instability, slow performance, non-responsive features, and repeated authentication failures. Account access issues appeared in $154,032$ reviews, while permissions and user information were mentioned in $38,126$ and $38,883$ reviews respectively. Confidentiality concerns, including explicit mentions of privacy violations and PHI exposure, occurred in $15,176$ reviews. The category distributions are summarised in Table III. The results show that usability failures in authentication and data access closely align with security weaknesses, indicating shared engineering flaws that undermine both stability and user trust.
@@ VI-A Platform Policy Enforcement
In this study, we examined Android applications, whose permission framework is designed for least-privilege use, yet enforcement is uneven. Many apps in our dataset bypass consent by exploiting undeclared or proprietary permissions, often accessing sensitive data such as location, SMS, and device identifiers. Google Play’s review process prioritizes malware detection but rarely flags static privacy violations, including insecure certificate configurations and privileged API use [muhammad2023circumventing, andow2019policylint]. To close this gap, marketplaces should embed automated tools such as RiskInDroid [merlo2017riskindroid], MobSF [bergadano2020modular], and OWASP Mobile Audit [acharya2015owasp] directly into submission workflows. Extending these checks to silent permission use, legacy cryptography, and telephony misuse, combined with manifest-to-behavior audits and update-time differential analysis, would help detect and prevent privilege escalation.
@@ VI-B By-Design Engineering
Weak enforcement is compounded by insecure development practices. Nearly half of evaluated apps still rely on deprecated cryptography (SHA-1, MD5), and known exploits such as Janus and StrandHogg 2.0 remain exploitable. This reflects inconsistent integration of SAST/DAST tools into CI/CD pipelines. Developers should adopt the OWASP MASVS [holguera2022owasp], embed security linters into IDEs, and align authentication with NIST SP 800-63B [chatzipavlou2016recommended]. To address risks from third-party SDKs, SBOM generation and automated dependency scanning should become prerequisites for app store submission [stoddard2023software].
Users notice when app behavior conflicts with stated purpose. Permission-function mismatches, especially for location, SMS, and contacts, frequently erode trust [lupton2012m]. Progressive, context-aware permission flows [bai2010context] can reduce suspicion, while platform rules should prohibit dark patterns that coerce consent, in line with EU Digital Services Act requirements [heldt2022eu]. At scale, NLP-based monitoring of app reviews could surface privacy concerns early, allowing developers to intervene before reputational damage occurs.
@@ VI-C Regulation of Gray-Zone Health Apps
Trust also depends on regulatory coverage. Frameworks such as HIPAA and GDPR focus on clinical providers, leaving many gray-zone apps (e.g., fitness trackers, symptom checkers) outside formal oversight despite their collection of PHI-equivalent data [istepanian2022mobile]. Regulators should broaden the definition of health data handlers to cover any entity processing biometric or medical information [carayon2018challenges]. Privacy nutrition labels in app stores and independent certification schemes (e.g., SOC 2–style for mHealth) would provide clearer accountability and verifiable standards [loser2014security].
@@ VI-D Automated Vetting and Socio-Technical Oversight
Finally, sustaining improvements requires moving beyond manual audits. Continuous vetting pipelines that combine static and dynamic analysis, taint tracking, permission mapping, and traffic anomaly detection can expose covert channels and privilege escalation paths [nurgalieva2020security]. Sandboxed monitoring of CPU, sensor, and network activity adds runtime assurance. For decision-makers, explainable dashboards built on tools like MobSF and RiskInDroid can translate technical findings into actionable insights [stoddard2023software], supporting coordinated socio-technical oversight.
@@ VIII Conclusion
IX Acknowledgment
We thank the Data Agency and Security (DAS) Lab at George Mason University (GMU) and the University of Denver (DU), where this study was conducted. We also thank Dhiman Goswami for helping to edit the final version of this paper. All opinions expressed are solely those of the authors.